# Using AI-Based Coding Assistants in Practice: State of Affairs, Perceptions, and Ways Forward

### v1

@@ V Discussion & Implications
V-A What Needs Our Focus Now
@@
The results of our study support the existing notion about the current state of affairs of using AI assistants in software development as favorable and rapidly developing. We see that people adopted the technology in their workflows and see the code provided by AI as usable and fairly accurate. However, in line with previous studies [36, 37, 38], we see that people are cautious regarding AI-generated code being insecure. We perceive this serious drawback as an opportunity for the community to research and develop solutions to provide users with more safe and secure technology.
@@
Section IV-C indicates that in terms of activities, writing tests and writing natural language artifacts are good candidates for improvement. The respondents indicated that they find these tasks to be somewhat unpleasant and express their desire to have them automated. The most popular stages where developers use AI assistance in these stages are: generating tests, generating data and resources for tests (e.g. inputs and outputs), summarizing code into docstring, generating comments at specific points in the code. Combining the established collaboration between developers and AI on these tasks with the willingness to delegate activities to AI, they warrant attention from the community for possible innovations in the respective fields.
@@
In the other studied activities, we can see some common patterns that are also important to take into account for improving the support of AI tools. In this regard, among the most promising steps in Implementing new features, Bug triaging and Refactoring are (a) the generation of the new necessary code and (b) the summarization of code for a better understanding of the context of the change. We suggest that studies of possible improvements in those directions would be important for developers and impactful for the field.
@@
Finally, while the option to chat with the assistant is especially crucial when implementing new features, it is present in all the other activities as well. This indicates the importance of not only the “technical” aspects of the models (e.g., generating quality code and summarizing the code correctly) but also UI and UX-related aspects, specifically conversational functionality of AI assistants [39, 25].
V-B What Needs to Change in the Future
@@
Results presented in Section IV-D indicate a way towards greatly improving the AI experience for users. While many different reasons for not using assistants can be seen in Table II, still the majority of the respondents said that they are likely or very likely to use AI assistants for all activities if the shortcomings were overcome. Some of the reasons are more technical, while others are more fundamental, which makes it difficult to address them equally. For example, Lack of need for AI assistance is a more general thing that gradually changes over time. However, some of the main drawbacks can be overcome and require long-term research commitments. We see the possible space of innovation threefold: system, integration, and education.
@@
One direction is the improvement of base systems of AI assistants. Addressing AI-generated output being inaccurate and the Lack of compliance with non-functional requirements is crucial, with 17.7% and 3.2% of all responses mentioning these issues, respectively. Previous research shows that developers spend almost 50% of coding time in interaction with the LLM, with 35% dedicated to double-checking suggestions, since the generated code is limited in meeting both functional and non-functional requirements [20]. Therefore, resolving accuracy issues would also affect the Time inefficiency of AI tooling usage, mentioned in 5.5% of all responses.
@@ VI Threats to Validity
Generalizability. Our results are based on a specific sample of people and might not generalize. In particular, a large part of the sample represents subscribers to JetBrains products, which can introduce some bias into the results. At the same time, our sample of 481 people is large for studies in our field [52], and it is diverse in terms of programming languages and types of developed software.
@@
VII Conclusion
In this paper, we set out to study specific ways in which developers use AI assistants in different stages of the software development life-cycle, as well as the reasons why they do not use them. Our results show that respondents use AI assistants in various activities and various stages within them, but unequally. Developers indicated Writing tests and Writing natural language artifacts as the least enjoyable activities that they would want to delegate to an AI assistant. Within different activities, developers tend to use assistants for generation and summarization, and employ them for finding places in the code and applying code more rarely. In terms of the reasons for not using AI assistants, our thematic analysis revealed 20 diverse issues, with the main ones being Lack of need for AI assistance, AI-generated output being inaccurate, Lack of trust, and Lack of understanding of context by AI assistant. We also highlight issues raised by the participants in regard to individual activities, which can inform further research.

### v2

@@ 5 Discussion & Implications
5.1 What Needs Our Focus Now
@@
The results of our study support the existing notion about the current state of affairs of using AI assistants in software development as favorable and rapidly developing [12, 15]. We see that people adopted the technology in their workflows and see the code provided by AI as usable and fairly accurate. However, in line with previous studies [30, 31, 32], we see that people are cautious regarding AI-generated code being insecure. We perceive this serious drawback as an opportunity for the community to research and develop solutions to provide users with more safe and secure technology.
@@
Our high-level comparison of activities (Section 4.3) indicates that in terms of activities, writing tests and writing natural language artifacts are good candidates for improvement. The comparison of responses to different questions allows us to conclude that respondents find these tasks to be somewhat unpleasant and express their desire to have them automated. Our detailed subsequent questions about individual stages highlight how specifically respondents use AI assistants within these activities: generating tests, generating data and resources for tests (e.g. inputs and outputs), summarizing code into docstring, generating comments at specific points in the code. While some of these tasks are well-established, others are less studied — for example, using AI assistants for generating separate resources for tests requires more attention from the research community. Combining the established collaboration between developers and AI on these tasks with the willingness to delegate activities to AI, they warrant attention from the community for possible innovations in the respective fields.
@@
In the other studied activities, we can see some common patterns that are also important to take into account for improving the support of AI tools. In this regard, among the most promising steps in Implementing new features, Bug triaging and Refactoring are (a) the generation of the new necessary code and (b) the summarization of code for a better understanding of the context of the change. These two activities align with the grounded theory proposed by Barke et al. [18], which describes two developers’ interaction modes with AI assistants — acceleration (supported by code generation) and exploration (supported by code summarization). Our detailed results for individual stages can highlight some more specific candidates for focus — Using code comprehension to summarize the most likely causes of the bug, Generating the refactored version of the code, and others. We suggest that studies of possible improvements in those directions would be important for developers and impactful for the field.
@@
Finally, while the option to chat with the assistant is especially crucial when implementing new features, it is present in all the other activities as well. This indicates the importance of not only the “technical” aspects of the models (e.g., generating quality code and summarizing the code correctly) but also UI and UX-related aspects, as mentioned by Zhou et al. [11], specifically conversational functionality of AI assistants as highlighted in works of Ross et al.[33] and Ziegler et al. [15].
These results can be formulated into the following general implications:
  • For researchers. To bring the most immediate value in the areas where developers already highlight their needs, one can focus on the innovation around the systems that ensure usage of AI assistants in writing tests, writing natural language artifacts, code generation, and code summarization.
  • For tool builders. The specific features that developers are already actively using — generating tests, summarizing code into docstrings, etc. — are crucial for better integration into the existing tools, such as IDEs. Some of the used features might not yet be directly available in the workflow, such as generating data for the tests, and it is worth experimenting with them.
5.2 What Needs to Change in the Future
@@
Results presented in Section 4.4 indicate a way towards greatly improving the AI experience for users. While many different reasons for not using assistants can be seen in Table 2, still the majority of the respondents said that they are likely or very likely to use AI assistants for all activities if the shortcomings were overcome. The large-scale nature of our survey provides both technical and more fundamental reasons, which makes it difficult to address them equally. For example, Lack of need for AI assistance is a more general thing that gradually changes over time. However, the wide array of answers from our respondents, referring to multiple different activities and stages within them, allows us to propose a multi-faceted look at the future research to tackle the existing shortcomings. We see the possible space of innovation threefold: system, integration, and education.
@@
One direction is the improvement of base systems of AI assistants. Addressing AI-generated output being inaccurate and the Lack of compliance with non-functional requirements is crucial, with 17.7% and 3.2% of all responses mentioning these issues, respectively, which is in line with the study of Liang et al. [12]. Previous research shows that developers spend almost 50% of coding time in interaction with the LLM, with 35% dedicated to double-checking suggestions, since the generated code is limited in meeting both functional and non-functional requirements [17]. Therefore, resolving accuracy issues would also affect the Time inefficiency of AI tooling usage, mentioned in 5.5% of all responses.
@@
These results can be formulated into the following general implications:
  • For researchers. With the recent advances in developing AI assistants, a lot of different problems remain, all of which require long-term research. In particular, it is important to focus not only on the systems themselves — addressing common issues like lack of quality, understanding larger context, and ensuring legal and security compliance — but also on the human element. Research should align the system with human needs, understanding the reasons behind certain challenges to effectively resolve them and create more user-centered solutions.
  • For tool builders. A good AI model in itself is not enough to bring value to the user. It is crucial to develop novel ways of integrating AI assistants into the developers’ workflow and invest in better UI/UX. With industrial experience in these kinds of improvements, this can be handled by the development teams of tools such as IDEs.
  • For developers. While the tools should improve their education of users on AI assistants, it is also on developers to inform themselves. Better knowledge of various features and experimenting with new ones can vastly improve the experience, as noted by some responses to our questions that highlighted the desire of the participants to try out new things that they picked up in our survey.
@@ 6 Threats to Validity
Generalizability. Our results are based on a specific sample of people and might not generalize. In particular, even though the mailing list curated by JetBrains includes users of different IDEs and AI assistants, it can introduce some bias into the results. At the same time, our sample of 481 people is large for studies in our field [47], and it is diverse in terms of programming languages and types of developed software.
@@
7 Conclusion
In this paper, we set out to study specific ways in which developers use AI assistants in different stages of the software development life-cycle, as well as the reasons why they do not use them. Our results show that respondents use AI assistants in various activities and various stages within them, but unequally. Developers indicated Writing tests and Writing natural language artifacts as the least enjoyable activities that they would want to delegate to an AI assistant. Within different activities, developers tend to use assistants for generation and summarization, and employ them for finding places in the code and applying code more rarely. In terms of the reasons for not using AI assistants, our thematic analysis revealed 20 diverse issues, with the main ones being Lack of need for AI assistance, AI-generated output being inaccurate, Lack of trust, and Lack of understanding of context by AI assistant. We also highlight issues raised by the participants in regard to individual activities, which can inform further research. The survey and its full results can be found in supplementary materials [13].
@@ 7 Conclusion
Acknowledgements
We would like to extend our sincere gratitude to the Surveys Team and the Research Core Team of JetBrains for their invaluable support and assistance throughout this study. Their expertise and dedication were crucial in facilitating the research process and ensuring its success. We greatly appreciate their contributions and collaboration.